# GeneMAN: Generalizable Single-Image 3D Human Reconstruction from Multi-Source Human Data

### v2

@@ D.3 Ablation of Multi-Source Human Dataset
[…]
@@
Table 5: Model Efficiency. Comparison of inference times between our model and the baselines. Note that PIFu [47], GTA [71], and SiTH [13] are feed-forward methods, whereas TeCH [18] and GeneMAN (ours) are optimization-based approaches that require per-subject optimization. Our GeneMAN model requires a total of 1.42 hours to generate a 3D human asset, with the following time breakdown: Geometry Initialization (22 min), Geometry Sculpting (15 min), Latent Space Texture Refinement (37 min), and Pixel Space Texture Refinement (11 min). All methods are tested on a single NVIDIA A100 80GB GPU.
PIFu GTA TeCH SiTH GeneMAN(Ours)
@@
Appendix E Limitations
Although our method achieves superior reconstruction performance on in-the-wild images compared to state-of-the-art approaches, it requires a longer optimization time to generate each 3D asset compared to feed-forward methods such as PIFu [47], GTA [71], and SiTH [13], which may limit its practical applicability. Nevertheless, our approach remains faster than TeCH [18], offering a $59.4\%$ increase in efficiency while delivering better quality. The model efficiency of both baseline methods and ours are reported in Tab. 5. Regarding reconstruction quality, our method still struggles to achieve fine-grained modeling for certain parts, such as the hands of full-body humans. Additionally, it lacks specific designs for handling occluded individuals. For human-with-objects reconstruction, our approach primarily focuses on reconstructing personal belongings, but it performs poorly with particularly large or complex objects, such as bicycles. Research on human-with-object reconstruction will be a key focus of our future work.
@@ Appendix E Limitations
Figure 16: More GeneMAN Results with Complex Poses on CAPE [33]. Best view with zoomed in.

### v3

@@ C.4 Ablation of Multi-Source Human Dataset
Methods PSNR $\uparrow$ LPIPS $\downarrow$ CLIP-Sim $\uparrow$
@@
Table 6: Model Efficiency. Comparison of inference times between our model and the baselines. Note that PIFu Saito et al. (2019), GTA Zhang et al. (2024b), and SiTH Ho et al. (2024) are feed-forward methods, whereas TeCH Huang et al. (2024b) and GeneMAN (ours) are optimization-based approaches that require per-subject optimization. Our GeneMAN model requires a total of 1.42 hours to generate a 3D human asset, with the following time breakdown: Geometry Initialization (22 min), Geometry Sculpting (15 min), Latent Space Texture Refinement (37 min), and Pixel Space Texture Refinement (11 min). All methods are tested on a single NVIDIA A100 80GB GPU.
PIFu Saito et al. (2019) GTA Zhang et al. (2024b) TeCH Huang et al. (2024b) SiTH Ho et al. (2024) GeneMAN(Ours)
@@
Appendix D Limitations
Although our method achieves superior reconstruction performance on in-the-wild images compared to state-of-the-art approaches, it requires a longer optimization time to generate each 3D asset compared to feed-forward methods such as PIFu Saito et al. (2019), GTA Zhang et al. (2024b), and SiTH Ho et al. (2024), which may limit its practical applicability. Nevertheless, our approach remains faster than TeCH Huang et al. (2024b), offering a $59.4\%$ increase in efficiency while delivering better quality. The model efficiency of both baseline methods and ours are reported in Tab. 6. Regarding reconstruction quality, our method still struggles to achieve fine-grained modeling for certain parts, such as the hands of full-body humans. Additionally, it lacks specific designs for handling occluded individuals. For human-with-objects reconstruction, our approach primarily focuses on reconstructing personal belongings, but it performs poorly with particularly large or complex objects, such as bicycles. Research on human-with-object reconstruction will be a key focus of our future work.
Appendix E Broader Impacts and Safeguards
GeneMAN introduces a generalizable framework for reconstructing high-quality 3D human models from a single in-the-wild image. This method is capable of handling various body proportions, poses, clothing, and personal belongings, offering a wide range of applications, including virtual reality (VR), augmented reality (AR), telepresence, digital human interfaces, film production, and 3D game development. However, the widespread application of GeneMAN may also pose risks. For instance, 3D human reconstruction technology could be misused to generate synthetic content, which may raise ethical and legal concerns. To ensure that the application of GeneMAN is positive and sustainable, we have implemented the following safeguards: 1) Ensure that the technological outcomes of GeneMAN are equitably accessible to diverse social groups, including marginalized communities. 2) Actively collaborate with communities and stakeholders to ensure that technological applications meet societal needs and expectations. 3) Ensure that the development and application of GeneMAN comply with all relevant laws and regulations, including data protection laws and intellectual property rights. 4) Respect all intellectual property rights and ensure that all used data and methods have been appropriately authorized.
@@ Appendix E Broader Impacts and Safeguards
Figure 16: More GeneMAN Results with Complex Poses on CAPE Ma et al. (2020). Best view with zoomed in.